Express V= 2t² + 5t + 9 as a linear combination of the polynomials

P1= t + 1

P2= t - 1

P3 = t² - 2t + 1


1
Expert's answer
2021-12-20T07:39:37-0500

Let


"V= 2t^2 + 5t + 9"

"=a(t+1)+b(t-1)+c(t^2- 2t + 1)"

Then


"2t^2 + 5t + 9=at+a+bt-b+ct^2-2ct+c"

"=ct^2+(a+b-2c)t+(a-b+c)"

"t^2:2=c"

"t^1:5=a+b-2c"

"t^0=9=a-b+c"

"c=2"

"2a-c=14"

"2b-3c=-4"

"a=8, b=1, c=2"

"V= 2t^2 + 5t + 9"

"=8(t+1)+1(t-1)+2(t^2- 2t + 1)"